# Approximation of multipartite quantum states: revised version with new applications

### v1

@@ 5.3 On the energy-constrained versions of $E_{R,\pi}$
where $H_{\!A^{n}}$ is the Hamiltonian of a $n$-partite system $A^{n}\doteq A_{1}...A_{n}$ [4]. [^17] In [4] the case when $E_{R,\pi}=E_{R}$ (i.e. when $\pi$ is the finest partition of $\{1,2,...,n\}$) is considered.
@@
where $H_{A_{s}}$ is a densely defined positive operator on $\mathcal{H}_{A_{s}}$, $s=\overline{1,n}$, then the set of $\pi$-separable states with finite energy is dense in $\mathfrak{S}^{\pi}_{\mathrm{sep}}(\mathcal{H}_{A_{1}...A_{n}})$. So, in this case it is reasonable to assume the coincidence of $E^{H_{\!A^{n}}}_{R,\pi}(\rho)$ and $E_{R,\pi}(\rho)$ for any state $\rho$ with finite energy. [^18] This coincidence is not obvious, since the infima of a lower semicontinuous function over a closed set and over a dense subset if this set may be different. In the following proposition we establish the validity of this assumption under the particular condition on the Hamiltonians $H_{A_{1}}$,…,$H_{A_{n}}$.
@@
$E^{H_{\!A^{n}}}_{R,\pi}(\rho)=E_{R,\pi}(\rho)$ (58)
for any state $\rho$ such that $\mathrm{Tr}H_{\!A^{n}}\rho=\sum_{s=1}^{n}\mathrm{Tr}H_{A_{s}}\rho_{A_{s}}<+\infty$. Moreover, for any such state $\rho$ the infimum in the definition (36) of $E_{R,\pi}(\rho)$ can be taken only over all finitely-decomposable separable states $\sigma$ (i.e. the states having the form (48)) such that $\mathrm{Tr}H_{\!A^{n}}\sigma=\sum_{s=1}^{n}\mathrm{Tr}H_{A_{s}}\sigma_{A_{s}}<+\infty$.
Proof. Assume that the Hamiltonians $H_{A_{1}}$,…,$H_{A_{n-1}}$ satisfy condition (10). Let $\mathfrak{S}_{0}$ be the convex subset of $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of all states $\rho$ with finite value of $\mathrm{Tr}H_{\!A^{n}}\rho=\sum_{s=1}^{n}\mathrm{Tr}H_{A_{s}}\rho$. Consider the function
@@
By the proof of Theorem 1 in [28] the set $\mathfrak{S}_{0}$ has the invariance property stated in Remark 9 in [31]. By this remark Theorem 5 in [31] is generalized to all the functions from the class $L_{n}^{n-1}(2,1|\hskip 1.0pt\mathfrak{S}_{0})$ containing the function $\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}$. Since the function $E_{R,\pi}$ belongs to the class $L_{n}^{n-1}(1,1)$, Theorem 5 in [31], its generalization mentioned before and the arguments from the proof of Theorem 3 in Section 3 show that
$\lim_{r\to+\infty}E_{R,\pi}(\rho_{r})=E_{R,\pi}(\rho)\quad\textrm{and}\quad\lim_{r\to+\infty}\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho_{r})=\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho),\quad\forall\rho\in\mathfrak{S}_{0},$ (59)
@@
where $\rho_{r}=c_{r}^{-1}Q_{r}\rho\hskip 1.0ptQ_{r}$, $\,Q_{r}=P_{r}^{1}\otimes...\otimes P_{r}^{n}$, $c_{r}=\mathrm{Tr}Q_{r}\rho$. Hence, to prove that $\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho)=E^{H_{\!A^{n}}}_{R,\pi}(\rho)=E_{R,\pi}(\rho)$ it suffices to show that $\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho_{r})=E_{R,\pi}(\rho_{r})$ for any $r$. This can be done by using Lemma 5.1 in Section 5.1 and by noting that $\mathrm{Tr}H_{\!A^{n}}\sigma=\sum_{s=1}^{n}\mathrm{Tr}H_{A_{s}}\sigma_{A_{s}}<+\infty$ for any $\pi$-separable state $\sigma$ supported by the subspace $P_{r}^{1}\otimes...\otimes P_{r}^{n}(\mathcal{H}_{A_{1}...A_{n}})$. $\square$
Appendix: On the classes $\widehat{L}^{m}_{n}(C,D)$ and $N^{m}_{n,s}(C,D)$
Below a description of the classes $\widehat{L}^{m}_{n}(C,D)$ and $N^{m}_{n,s}(C,D)$ of functions on the set of states of a composite quantum system $A_{1}...A_{n}$ is presented along with a list of basic characteristics belonging to these classes. The functions in $\widehat{L}^{m}_{n}(C,D)\cup N^{m}_{n,s}(C,D)$ has a faithful common $II^{m}_{n}$-type continuity bound under the energy-type constraints in the subsystems $A_{1}$,..,$A_{m}$ given by expression (19). [^19] If $A_{1}$,…,$A_{m}$ are finite-dimensional quantum systems then a faithful common Fannes’ type continuity bound for the functions in $\widehat{L}^{m}_{n}(C,D)\cup N^{m}_{n,s}(C,D)$ is given by Theorem 3 in [33, Section 3].
Let $m\leq n$ and $L^{m}_{n}(C,D)$ be the class of all (finite) functions $f$ on the set [^20] We assume that $A_{1}$,…,$A_{n}$ are infinite-dimensional quantum systems.

### v2

@@ 4.1.5 On the energy-constrained versions of $E_{R,\pi}$
where $H_{\!A^{n}}$ is the Hamiltonian of $n$-partite quantum system $A^{n}\doteq A_{1}...A_{n}$ [55]. [^21] In [55] the case when $E_{R,\pi}=E_{R}$ (i.e. when $\pi$ consists of the single finest partition of $\{1,2,...,n\}$) is considered.
@@
where $H_{A_{s}}$ is a positive operator on $\mathcal{H}_{A_{s}}$, $s=\overline{1,n}$, then the set of $\pi$-separable states with finite energy is dense in $\mathfrak{S}^{\pi}_{\mathrm{sep}}(\mathcal{H}_{A_{1}...A_{n}})$. So, in this case it is reasonable to assume the coincidence of $E^{H_{\!A^{n}}}_{R,\pi}(\rho)$ and $E_{R,\pi}(\rho)$ for any state $\rho$ with finite energy. [^22] This coincidence is not obvious, since the infima of a lower semicontinuous function over a closed set and over a dense subset if this set may be different. In the following proposition we establish the validity of this assumption under the particular condition on the Hamiltonians $H_{A_{1}}$,…,$H_{A_{n}}$.
@@
$E^{H_{\!A^{n}}}_{R,\pi}(\rho)=E_{R,\pi}(\rho)$
for any state $\rho$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ such that $\mathrm{Tr}H_{\!A^{n}}\rho=\sum_{s=1}^{n}\mathrm{Tr}H_{A_{s}}\rho_{A_{s}}<+\infty$. Moreover, for any such state $\rho$ the infimum in the definition (64) of $E_{R,\pi}(\rho)$ can be taken only over all finitely-decomposable $\pi$-separable states $\sigma$ (i.e. the states having the form (69)) such that $\mathrm{Tr}H_{\!A^{n}}\sigma=\sum_{s=1}^{n}\mathrm{Tr}H_{A_{s}}\sigma_{A_{s}}<+\infty$.
Proof. By the arguments in Remark 4.1.1 we may assume that the Hamiltonians $H_{A_{1}}$,…,$H_{A_{n-1}}$ satisfy condition (17). We may also assume that all the Hamiltonians $H_{A_{1}}$,…,$H_{A_{n}}$ are densely defined operators.
Let $\mathfrak{S}_{0}$ be the convex subset of $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ consisting of all states $\rho$ with finite value of $\mathrm{Tr}H_{\!A^{n}}\rho=\sum_{s=1}^{n}\mathrm{Tr}H_{A_{s}}\rho$. Consider the function
@@
By the proof of Theorem 1 in [20] the set $\mathfrak{S}_{0}$ has the invariance property stated in Remark 9 in [15]. By this remark Theorem 5 in [15] is generalized to all the functions from the class $L_{n}^{n-1}(2,1|\hskip 1.0pt\mathfrak{S}_{0})$ containing the function $\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}$. Note also that the function $E_{R,\pi}$ belongs to the class $L_{n}^{n-1}(1,1)$ [6, the proof of Proposition 39]. Thus, Theorem 5 in [15], its generalization mentioned before and the arguments from the proof of Theorem 3.1 in Section 3 show that
$\lim_{r\to+\infty}E_{R,\pi}(\rho_{r})=E_{R,\pi}(\rho)\quad\textrm{and}\quad\lim_{r\to+\infty}\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho_{r})=\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho),\quad\forall\rho\in\mathfrak{S}_{0},$
@@
where $\rho_{r}=c_{r}^{-1}Q_{r}\rho\hskip 1.0ptQ_{r}$, $\,Q_{r}=P_{r}^{1}\otimes...\otimes P_{r}^{n}$, $c_{r}=\mathrm{Tr}Q_{r}\rho$. Hence, to prove that $\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho)=E^{H_{\!A^{n}}}_{R,\pi}(\rho)=E_{R,\pi}(\rho)$ it suffices to show that $\widetilde{E}^{H_{\!A^{n}}}_{R,\pi}(\rho_{r})=E_{R,\pi}(\rho_{r})$ for any $r$. This can be done by using Lemma 4.1.2 in Section 4.1.2 and by noting that $\mathrm{Tr}H_{\!A^{n}}\sigma=\sum_{s=1}^{n}\mathrm{Tr}H_{A_{s}}\sigma_{A_{s}}<+\infty$ for any $\pi$-separable state $\sigma$ in $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$ supported by the subspace $P_{r}^{1}\otimes...\otimes P_{r}^{n}(\mathcal{H}_{A_{1}...A_{n}})$. $\square$
4.2 Rains bound and its regularization
4.2.1 Definitions and basic properties
The Rains bound (also called Rains Divergence) is an important characteristic of a bipartite quantum state that gives an upper bound on the distillable entanglement of this state [50, 56, 57]. By using the Lindblad’s extension of the quantum relative (defined in (6)) the Rains bound can be defined as
$R(\rho)=\inf_{\sigma\in\mathfrak{T}_{R}}(D(\rho\|\sigma)+1-\mathrm{Tr}\sigma),$
where
$\mathfrak{T}_{R}=\left\{\sigma\in\mathfrak{T}(\mathcal{H}_{AB})\,|\,\sigma\geq 0,\|\sigma^{\Gamma}\|_{1}\leq 1\right\},$ (85)
$\Gamma$ is the partial transposition map in $\mathfrak{T}(\mathcal{H}_{AB})$ [58]. The Rains bound possesses many properties of entanglement measures, in particular, it is a convex function equal to zero on the set of separable states and non-increasing under LOCC operations. However, the Rains bound may be equal to zero at some entangled states (f.i., it is equal to zero at all the entangled states with positive partial transpose). The main advantage of the Rains bound as a quasi-entanglement measure consists in the fact that it gives a tighter upper bound on the distillable entanglement than the relative entropy of entanglement. It is essential also that it has a semi-definite program formulation [56, 58].
Since the set $\mathfrak{T}_{R}$ (defined in (85)) contains all separable states in $\mathfrak{S}(\mathcal{H}_{AB})$, the Rains bound $R(\rho)$ does not exceed the relative entropy of entanglement $E_{R}$:
$R(\rho)\leq E_{R}(\rho)\quad\forall\rho\in\mathfrak{S}(\mathcal{H}_{AB}).$
4.2.2 Continuity and convexity properties of $R$ and $R^{\infty}$
Despite the fact that the Rains bound definition is very similar to the definition of the relative entropy of entanglement, the technique developed in [6] for analysis of the latter function (and other ”relative entropies of resource”) in the infinite-dimensional case can not be applied to the Rains bound (by the reason described at the end of Section 4.3 in [6]). So, the good analytical properties (achievability and lower semicontinuity) of the function $E_{R}$ and of the relative entropy of PPT-entanglement in infinite-dimensional bipartite systems established in [6] are not yet proved for the Rains bound. Accordingly, the convergence criterion and all its corollaries obtained in [7] can not also be applied to the Rains bound. So, one can say that for all its advantages, the Rains bound is more difficult for analysis in the infinite-dimensional case than other entanglement measures of the relative entropy distance type.
The approximation technique developed in Section 3 gives a way to partially overcome the problems with analysis of the Rains bound mentioned before. It also allows us to establish analytical properties for the regularized Rains bound defined as
$R^{\infty}(\rho)=\lim_{k\rightarrow+\infty}k^{-1}R(\rho^{\otimes k})=\inf_{k\in\mathbb{N}}k^{-1}R(\rho^{\otimes k}),$
where $\rho^{\otimes k}$ is treated as a state of the bipartite quantum system $(A^{k})(B^{k})$. The above limit exists due to Fekete’s lemma (cf.[47]), because the Rains bound is subadditive on tensor products. The nonadditivity of the Rains bound implies that $R^{\infty}\neq R$ [59].
The results in Section 3 can be applied to the functions $R$ and $R^{\infty}$ because both these functions belong to the class $II_{2}^{1}$: a common faithful $II_{2}^{1}$-type continuity bound for $R$ and $R^{\infty}$ is presented in Proposition 36 in [6].
Proposition 9.  Let